log(1/10^-3) = 3
Thats the answer I get with the calculator but whats the step by step to reach this answer?

Answer by lwsshak3(11628)

log(1/10^-3) = 3
Thats the answer I get with the calculator but whats the step by step to reach this answer?
log(1/10^-3) = 3
log(10^3)=3
rule: log of base raised to a power=power
in this example:
base=10
log(10^3)=3